OpenStudy (anonymous):

Indices. If 2^x=m and 3^x=n, 8^(x+1) + 4(27^(1/3-x)). express the following in terms of m and n.

OpenStudy (anonymous):

8*m^3+12/(n^3)

OpenStudy (anonymous):

Can you show me the working please?

OpenStudy (anonymous):

8*8^x+4(3/27^3)=8*2^3x+12/3^3x

OpenStudy (anonymous):

2^3x=m^3 3^3x=n^3